Calculate the force exerted on a car of mass 1 ton, that moved with velocity of 45km/hr on a rough road and it is brought to rest after 10 seconds
1
Expert's answer
2019-08-02T10:26:32-0400

The Newton's second law states

"F=ma"

The acceleration of a car

"a=\frac{\Delta v}{\Delta t}=\frac{0\:\rm{m/s}-45(\rm{1000\:m/3600\: s})}{10\:\rm{s}}=-1.25\:\rm{m/s^2}"

Finally, the magnitude of the force exerted on a car

"F=1000\:\rm{kg}\times 1.25\:\rm{m/s^2}=1250\:\rm{N}"